Find the amount of interest and the monthly payment for the loan $3,400 at 12%
add-on interest for 2 years.

Answer by Theo(13342)

12% per year compounded monthly results in 1% per month (12/12 = 1).
the loan amount is 3400.
the period of the loan is 24 months.
the monthly payment is 160.0498056
over the 24 month period of the loan, you will have made a total payments of 24 * 160.0498056 = 3841.195333.
the interest you paid on the loan is 3841.195333 - 3400 = 441.195333.
i used a financial calculator to provide your answer.
i'm not sure what you're using.
let me know if this is enough for what you need to know or you need more details on how it was done.
let me know what tools you have available to you so i can tailor the answer to that.